# UrbanKGent: A Unified Large Language Model Agent Framework for Urban Knowledge Graph Construction

### v1

@@ Abstract.
Urban knowledge graph has recently worked as an emerging building block to distill critical knowledge from multi-sourced urban data for diverse urban application scenarios. Despite its promising benefits, urban knowledge graph construction (UrbanKGC) still heavily relies on manual effort, hindering its potential advancement. This paper presents UrbanKGent, a unified large language model agent framework, for urban knowledge graph construction. Specifically, we first construct the knowledgeable instruction set for UrbanKGC tasks (such as relational triplet extraction and knowledge graph completion) via heterogeneity-aware and geospatial-infused instruction generation. Moreover, we propose a tool-augmented iterative trajectory refinement module to enhance and refine the trajectories distilled from GPT-4. Through hybrid instruction fine-tuning with augmented trajectories on Llama-2-13B, we obtain the UrbanKGC agent, UrbanKGent-13B. We perform a comprehensive evaluation on two real-world datasets using both human and GPT-4 self-evaluation. The experimental results demonstrate that UrbanKGent-13B not only can significantly outperform 21 baselines in UrbanKGC tasks, but also surpass the state-of-the-art LLM, GPT-4, by more than 10% with approximately 20 times lower cost. We deploy UrbanKGent-13B to provide online services, which can construct an UrbanKG with thousands of times richer relationships using only one-fifth of the data compared with the existing benchmark. Our data, code, and opensource UrbanKGC agent are available at https://github.com/usail-hkust/UrbanKGent. Refer to caption Figure 1. The framework overview of UrbanKGent.
urban knowledge graph, knowledge graph construction, large language model agent
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligence; Natural language processing; Information extraction
1. Introduction
@@
Urban Knowledge Graph (UrbanKG) aims to model intricate relationships and semantics within a city by extracting and organizing urban entities (e.g., POIs, road networks, etc.) into a multi-relational heterogeneous graph. As an emerging building block, multi-sourced urban data are widely used to construct an UrbanKG to provide critical knowledge for various knowledge-enhanced urban downstream tasks, such as traffic management, pollution monitoring, and emergency response (Sun et al., 2017; Chen et al., 2016; Cheng et al., 2018). UrbanKG has gradually become an essential tool of the modern smart city.
@@
In prior literature, many efforts have been devoted to urban knowledge graph construction (UrbanKGC) using massive urban data sources. In particular, one commonly used approach (Wang et al., 2021; Liu et al., 2022; Ning et al., 2023) is to extract entities from structured urban data (e.g., geographic data, city sensor data, and traffic data) and define the relationships between obtained urban entities based on manually designed rules or patterns. However, these approaches suffer heavy reliance on a deep understanding of the application domain and are labor-intensive. Recently, inspired by the success of the Large Language Models (LLMs) in knowledge graph construction (Xie et al., 2023; Zheng et al., 2021; Yao et al., 2023), the LLMs have been adopted to improve UrbanKGC. For instance, GeoLM (Wang et al., 2023b) is proposed for spatial grounded entity recognition and relation extraction. K2 (Deng et al., 2023) retrains the Llama-2-7B model on a huge annotated geoscience text corpus for geospatial relation extraction. Nevertheless, these works rely on annotated corpus and model retraining, which may discourage researchers from adopting it for their own work and thus limit the application of UrbanKG.
LLM agent (et al, 2023a, b) has recently emerged and shown remarkable zero-shot capability for autonomous domain-specific task completion. For example, Voyager (Wang et al., 2023c) is a LLM-powered agent for zero-shot game exploration without re-training, and LLMLight (Gur et al., 2023) is a traffic signal control agent with zero-shot LLM reasoning ability. These studies motivate us to construct tailored LLM agents to address the aforementioned limitations in UrbanKG construction.
@@
Figure 2. Illustrative example of urban relational triplet extraction and knowledge graph completion. (a) The heterogeneous relationship understanding limitation of LLMs can be addressed by injecting prior urban knowledge into instruction. (b) The geospatial computing limitation of LLMs can be alleviated by invoking external geospatial tools.
@@
In fact, constructing an LLM agent compatible with various UrbanKGC tasks is a non-trivial problem due to the following two challenges: (1) Challenges 1: How to adapt LLMs for UrbanKGC? LLMs may not align well with the specific task due to the gap (Fu et al., 2023b) between the natural language processing corpus for training LLMs and the domain-specific corpus in urban domain (Lin et al., 2023). For example, the urban text data is usually heterogeneous and contains multifaceted urban knowledge (e.g., spatial, temporal, and functional aspects) (Deng et al., 2023). As shown in Figure 2(a), the text description of ”Columbia University” reflects its geographic spatial locations (i.e., spatial relationship), construction timelines (i.e., temporal relationship), and how it provides educational service for the city (i.e., functional relationship). LLMs may require a tailored alignment to understand heterogeneous urban relationships to extract these urban spatial, temporal, and functional relations accurately. (2) Challenges 2: How to improve the capacity of LLMs for UrbanKGC? The effectiveness of LLMs for urban knowledge graph construction is restricted by their feeble numerical computation capacity (Frieder et al., 2023; Zhang et al., 2023b), leading to their disability in complex geospatial relationship extraction (Bhandari et al., 2023; Mooney et al., 2023). However, the urban geospatial relationship plays a vital role in urban semantic modeling (Li et al., 2023c) and has been widely incorporated in previous UrbanKGs (Ning et al., 2023; Liu et al., 2023). As can be seen in Figure 2, extracting ”disconnected” relation between the geo-entity ”Columbia University” and ”Empire State Building” is useful for urban geo-semantic modeling. Accurately extracting such geospatial relationship demands necessary geospatial computing (e.g., utilizing latitude and longitude for distance calculation) and reasoning (i.e., deriving calculation results for geospatial relation reasoning) capabilities. It is appealing to improve the geospatial computing and reasoning ability (e.g., invoking external tools for calculation) of LLMs to satisfy the UrbanKGC task requirement.
@@
To address the aforementioned challenges, in this study, we propose UrbanKGent, a unified LLM agent framework for automatic UrbanKG construction. Figure 1 illustrates the overview of UrbanKGent. For a given city, we first generate a knowledgeable instruction set for UrbanKGC tasks (e.g., relational triplet extraction and knowledge graph completion) from urban geographic and text data sources. By heterogeneity-aware and geospatial-infused instruction generation, as shown in Figure 2(a), various urban spatiotemporal relationship knowledge can be encoded into the instruction, which facilitates alignment between LLMs with the target UrbanKGC tasks. Moreover, we propose a tool-augmented iterative trajectory refinement module to enhance and refine the trajectory derived by distilling GPT-4 with the above constructed instructions. Based on geospatial tool augmentation and self-refinement, the deficiency of LLMs in geospatial computing and reasoning could be alleviated, and unfaithful trajectories could be filtered out. Finally, we perform hybrid instruction fine-tuning based on the enhanced and refined trajectories on Llama 2-7B and Llama 2-13B variants (Touvron et al., 2023) by using LoRA (Hu et al., 2021). The obtained agent, UrbanKGent-13B, is feasible for completing multiple UrbanKGC tasks in a cost-effective way (i.e., no extra GPT-API cost needed).
@@
We conduct comprehensive experiments on two UrbanKGC tasks in two metropolises (New York City and Chicago) using both human evaluation and GPT-4-based self-evaluation. The empirical results validate the effectiveness of the proposed LLM agent for completing various UrbanKGC tasks. Moreover, we deploy the constructed agent, UrbanKGent-13B, online for UrbanKGC service, which can extract the same scale of triplets and entities of existing UrbanKG benchmark (Ning et al., 2023) using only one-fifth of data, and even expand the types of relations by thousands of times.
Our contributions are summarized as follows:
  • We propose the first UrbanKGC agent construction framework UrbanKGent, and release the UrbanKGent-13B to provide real-world UrbanKGC service [^1] https://htmlpreview.github.io/?https://raw.githubusercontent.com/usail-hkust/UrbanKGent/main/UrbanKGent%20Demo/index.html, which offers new opportunities to advance UrbanKG studies.
@@
  • We propose a knowledgeable instruction generation module and a tool-augmented iterative trajectory refinement method, which align LLMs to UrbanKGC tasks and compensate for their geospatial computing and reasoning inability.
  • Extensive experiments on two real-world datasets validate the effectiveness of the proposed framework and uncover its exceptional performance across various UrbanKGC tasks.
2. UrbanKGC Data Description
This section introduces the process of data collection and preprocessing details.
2.1. Data Collection
@@ 2.1.1. Geographic Data
The geographic data provides critical urban spatial structure information and functional semantics, which has been widely used in previous UrbanKG studies (Ning et al., 2023; Liu et al., 2023, 2021; Zhang et al., 2023a).
@@
Area-Of-Interst (AOI) Data. AOI data describes the urban spatial area structure, including urban commercial areas (e.g., shopping centers), residential areas (e.g., communities and neighbors), and so on. In this work, we first follow UUKG (Ning et al., 2023) to acquire the AOI name and geometry value from NYC Gov [^2] https://www.nyc.gov/ and CHI Gov [^3] https://www.chicago.gov/. Next, we use the AOI name to search for their text description from Wikipedia website [^4] https://www.wikipedia.org/. If there is no direct match, we then retrieve the Wikipedia pages that mention the AOI name and collect the related paragraphs. Each AOI record contains an AOI name, a polygon geometry value, and a text description. For example, [”Jamaica Bay”, polygon (-73.86 40.58, …), ”Jamaica Bay is an estuary …”] is the record of the AOI ”Jamaica Bay” with geometry value and text description.
@@
Road Network Data. Road data describes the urban spatial transportation network, including urban motorways, overpasses, and so on. We first follow (Ning et al., 2023) to obtain the road name and geometry value from Open Street Map (OSM) [^5] https://www.openstreetmap.org/. Then, following the same text acquisition operation in AOI data, we crawl the textual description of each road record from Wikipedia. Each road record contains a road name, a linestring geometry value, a road type and a text description. For example, [”Central Park Avenue”, linestring (-73.87 40.90, …), primary, ”Central Park Avenue is a boulevard in …”] describes the primary road named ”Central Park Avenue” with a linestring geometry value and its textual description.
@@
Point-Of-Interest (POI) Data. POI data represents different urban functions (e.g., residential and commercial), which have been widely adopted in many recent UrbanKG works (Liu et al., 2023; Wang et al., 2021; Ning et al., 2023). We first follow (Ning et al., 2023) to obtain the POI name, and geometry value from OSM. Then the textual description of each POI record could be crawled from Wikipedia following the similar process. Each POI record contains a POI name, a coordinate geometry, a POI type, and a text description. For example, [”Trump World Tower”, coordinate (-73.96 40.75), residential, ”Trump World Tower is a residential condominium …”] is the record of the POI ”Trump World Tower”.
2.1.2. Text Data
The text data provides rich contextual knowledge of the city space from different perspectives (e.g., the spatial context) (Deng et al., 2023), and it plays an important role in geospatial understanding. In this work, we collect two types of text corpus.
Review Data. The review of urban places provides useful commercial information that citizens use to make business decisions (Cho et al., 2011), which plays a critical role in urban knowledge distillation. We collect review data from Google Map [^6] https://www.google.com/maps. Specifically, we first manually split the city into multiple rectangular regions, then we utilize the Google Map API to query the places contained within each region and their reviews. Each review record contains a place name, a coordinate geometry value, a rating, and a text review. For example, [”Lifestyles Academy Inc”, coordinate (-87.87 41.65), 4.9, ”Very nice organization and …”] is the review record of place ”Lifestyles Academy Inc”.
Web Page Data. The web page data works as the general text corpus for the city, and it contains rich geoscience knowledge that has been utilized in recent urban entity and relation extraction studies (Deng et al., 2023). We collect web page data from the Google search engine. Specifically, we first input the name of the crawled AOI, Road, and POI record into Google. Then we concatenate the textual sentences of the top 10 retrieved web pages. Each web page record contains a long urban text description.
Table 1. The statistics of raw datasets.
@@ 2.1.2. Text Data
Geographic Data # of AOI 183 80
# of road 6,397 1,893
# of POI 5,369 5,658
@@
2.2. Data Preprocessing
Before constructing the UrbanKGC dataset, we first preprocess the raw datasets. Specifically, we filter out AOIs, roads, POIs, reviews, and web pages whose crawled textual descriptions are null value, too short (e.g., less than ten word description) or meaningless (e.g., just repeating the POI name). In addition, we remove irrelevant information from the text description, such as non-English characters, non-ASCII gibberish, website addresses, and so on. More details can be found in Appendix A.
3. Preliminary
@@
3.1. Task Definition and Problem Formulation
Before diving into the technical details, we first introduce the definition of urban knowledge graph (UrbanKG):
Definition 1 ().
@@ 3.1. Task Definition and Problem Formulation
The UrbanKG encodes diverse urban semantic knowledge by connecting urban entities into a multi-relational graph. This work aims to construct an UrbanKG from collected unstructured text data. We decompose the UrbanKG construction (UrbanKGC) process into two sequential knowledge graph construction tasks, namely relational triple extraction (Zheng et al., 2021) and knowledge graph completion (Yao et al., 2023). We first provide the basic definition for these two subtasks, and then introduce the problem formulation of this work.
3.1.1. Task Definition
We give the basic task definition as follow.

### v2

@@ Abstract
Urban knowledge graph has recently worked as an emerging building block to distill critical knowledge from multi-sourced urban data for diverse urban application scenarios. Despite its promising benefits, urban knowledge graph construction (UrbanKGC) still heavily relies on manual effort, hindering its potential advancement. This paper presents UrbanKGent, a unified large language model agent framework, for urban knowledge graph construction. Specifically, we first construct the knowledgeable instruction set for UrbanKGC tasks (such as relational triplet extraction and knowledge graph completion) via heterogeneity-aware and geospatial-infused instruction generation. Moreover, we propose a tool-augmented iterative trajectory refinement module to enhance and refine the trajectories distilled from GPT-4. Through hybrid instruction fine-tuning with augmented trajectories on Llama 2 and Llama 3 family, we obtain UrbanKGC agent family [^1] https://huggingface.co/usail-hkust/UrbanKGent-7B, https://huggingface.co/usail-hkust/UrbanKGent-8B and https://huggingface.co/usail-hkust/UrbanKGent-13B, consisting of UrbanKGent-7/8/13B version. We perform a comprehensive evaluation on two real-world datasets using both human and GPT-4 self-evaluation. The experimental results demonstrate that UrbanKGent family can not only significantly outperform 31 baselines in UrbanKGC tasks, but also surpass the state-of-the-art LLM, GPT-4, by more than 10% with approximately 20 times lower cost. Compared with the existing benchmark, the UrbanKGent family could help construct an UrbanKG with hundreds of times richer relationships using only one-fifth of the data. Our data and code are available at https://github.com/usail-hkust/UrbanKGent.
1 Introduction
@@
Urban Knowledge Graph (UrbanKG) aims to model intricate relationships and semantics within a city by extracting and organizing urban entities (e.g., POIs, road networks, etc.) into a multi-relational heterogeneous graph [1]. As an emerging building block, multi-sourced urban data are widely used to construct an UrbanKG to provide critical knowledge for various knowledge-enhanced urban downstream tasks, such as traffic management, pollution monitoring, and emergency response [2, 3, 4, 5]. UrbanKG has gradually become an essential tool of the modern smart city.
@@
In prior literature, many efforts have been devoted to urban knowledge graph construction (UrbanKGC) using massive urban data sources. In particular, one commonly used approach [6, 7, 8] is to extract entities from structured urban data (e.g., geographic data, city sensor data, and traffic data) and define the relationships between obtained urban entities based on manually designed rules or patterns. However, these approaches suffer heavy reliance on a deep understanding of the application domain and are labor-intensive. Recently, inspired by the success of the Large Language Models (LLMs)
@@
Figure 1: Illustrative example of urban relational triplet extraction and knowledge graph completion. (a) The heterogeneous relationship understanding limitation of LLMs can be addressed by injecting prior urban knowledge into instruction. (b) The geospatial computing limitation of LLMs can be alleviated by invoking external geospatial tools.
in knowledge graph construction [9, 10, 11], the LLMs have been adopted to improve UrbanKGC. For instance, GeoLM [12] is pre-trained on crowdsourced geographical corpus for geospatial entity recognition and relation extraction. K2 [13] retrains Llama-2-7B model on manually processed and filtered geoscience text corpus for geospatial relation extraction. Nevertheless, these works rely on high quality corpus annotation and computational extensive model retraining, which may discourage researchers from adopting UrbanKG for their own work.
LLM agent [14, 15] has recently emerged and shown remarkable zero-shot capability for autonomous domain-specific task completion. For example, Voyager [16] is a LLM-powered agent for zero-shot game exploration without re-training, and LLMLight [17] is a traffic signal control agent with zero-shot LLM reasoning ability. These studies motivate us to construct tailored LLM agents to address the aforementioned limitations in UrbanKG construction.
@@
In fact, constructing an LLM agent compatible with various UrbanKGC tasks is a non-trivial problem due to the following two challenges: (1) Challenges 1: How to adapt LLMs for UrbanKGC? LLMs may not align well with the specific task due to the gap [18] between the natural language processing corpus for training LLMs and the domain-specific corpus in urban domain [19]. For example, the urban text data is usually heterogeneous and contains multifaceted urban knowledge (e.g., spatial, temporal, and functional aspects) [13]. As shown in Figure 1(a), the text description of "Columbia University" reflects its geographic spatial locations (i.e., spatial relationship), construction timelines (i.e., temporal relationship), and how it provides educational service for the city (i.e., functional relationship). LLMs may require a tailored alignment to understand heterogeneous urban relationships to extract these urban spatial, temporal, and functional relations accurately. (2) Challenges 2: How to improve the capacity of LLMs for UrbanKGC? The effectiveness of LLMs for urban knowledge graph construction is restricted by their feeble numerical computation capacity [20, 21], leading to their disability in complex geospatial relationship extraction [22, 23]. However, the urban geospatial relationship plays a vital role in urban semantic modeling [12] and has been widely incorporated in previous UrbanKGs [8, 24]. As can be seen in Figure 1, extracting "disconnected" relation between the geo-entity "Columbia University" and "Empire State Building" is useful for urban geo-semantic modeling. Accurately extracting such geospatial relationship demands necessary geospatial computing (e.g., utilizing latitude and longitude for distance calculation) and reasoning (i.e., deriving calculation results for geospatial relation reasoning) capabilities. It is appealing to improve the geospatial computing and reasoning ability of LLMs to satisfy the UrbanKGC task requirement.
@@
To address the aforementioned challenges, in this study, we propose UrbanKGent, a unified LLM agent framework for automatic UrbanKG construction. Figure 2 illustrates the overview of UrbanKGent. For a given city, we first generate a knowledgeable instruction set for UrbanKGC tasks (relational triplet extraction and knowledge graph completion) from urban geographic and text data sources. By heterogeneity-aware and geospatial-infused instruction generation, as shown in Figure 1(a), various urban spatiotemporal relationship knowledge can be encoded into the instruction, which facilitates alignment between LLMs with the target UrbanKGC tasks. Moreover, we propose a tool-augmented iterative trajectory refinement module to enhance and refine the trajectory derived by distilling GPT-4 with the above constructed instructions. Based on geospatial tool augmentation and self-refinement, the deficiency of LLMs in geospatial computing and reasoning could be alleviated, and unfaithful trajectories could be filtered out. Finally, we perform hybrid instruction fine-tuning based on the enhanced and refined trajectories on Llama 2-7/13B and Llama 3-8B variants [25] by using LoRA [26]. The obtained UrbanKGent agent including 7/8/13B version, is feasible for completing multiple UrbanKGC tasks cost-effectively without extra GPT-API cost.
Figure 2: The framework of UrbanKGent.
@@
We conduct comprehensive experiments on two UrbanKGC tasks in two metropolises (New York City and Chicago) using both human evaluation and GPT-4-based self-evaluation. The empirical results validate the effectiveness of the proposed LLM agent for completing various UrbanKGC tasks. Moreover, the obtained UrbanKGent family (7/8/13B version) could help extract the same scale of triplets and entities of existing UrbanKG benchmark [8] using only one-fifth of data, and even expand the types of relations by hundreds of times.
@@
Our contributions are summarized as follows: (1) We propose the first UrbanKGC agent framework UrbanKGent and UrbanKGent family to provide real-world UrbanKGC service, offering new opportunities to advance UrbanKG studies. (2) We propose a knowledgeable instruction generation module and a tool-augmented iterative trajectory refinement method, which align LLMs to UrbanKGC tasks and compensate for their geospatial computing and reasoning inability. (3) Extensive experiments on two real-world datasets validate the effectiveness of proposed framework and uncover its exceptional performance across UrbanKGC tasks.
2 UrbanKGC Data Description
2.1 Data Collection
@@ 2.1.1 Geographic Data
The geographic data provides critical urban spatial structure information and functional semantics, which has been widely used in previous UrbanKG studies [8, 24, 27, 28].
@@
Area-Of-Interst (AOI) Data. AOI data describes the urban spatial area structure, including urban commercial areas (e.g., shopping centers), residential areas (e.g., communities), and so on. In this work, we first follow UUKG [8] to acquire the AOI name and geometry value from NYC Gov  [^2] https://www.nyc.gov/ and CHI Gov  [^3] https://www.chicago.gov/. Next, we use the AOI name to search their text description from Wikipedia and C4  [^4] https://huggingface.co/datasets/allenai/c4 dataset. Each AOI record contains an AOI name, a polygon geometry value, and a text description. For example, ["Jamaica Bay", polygon (-73.86 40.58, …), "Jamaica Bay is an estuary …"] is the record of the AOI "Jamaica Bay" with geometry value and text description.
@@
Road Network Data. Road data describes the urban spatial network, including urban motorways, overpasses, and so on. We first follow [8] to obtain the road name and geometry value from Open Street Map (OSM 555(https://www.openstreetmap.org/. Then, following the same text acquisition operation in AOI data, we crawl the textual description of each road record from Wikipedia. Each road record contains a road name, a linestring geometry value, a road type and a text description. For example, ["Central Park Avenue", linestring (-73.87 40.90, …), primary, "Central Park Avenue is a boulevard in …"] describes the primary road named "Central Park Avenue" with a linestring geometry value and its textual description.
@@
Point-Of-Interest (POI) Data. POI data represents different urban functions (e.g., residential and commercial), which have been widely adopted in many recent UrbanKG works [24, 6, 8]. We first follow [8] to obtain the POI name, and geometry value from OSM. Then the textual description of each POI record could be crawled from Wikipedia following the similar process. Each POI record contains a POI name, a coordinate geometry, a POI type, and a text description. For example,
Table 1: The statistics of raw datasets.
@@
Geographic Data # of AOI 192 136
# of road 6,765 2,241
# of POI 5,872 5,877
@@
["Trump World Tower", coordinate (-73.96 40.75), residential, "Trump World Tower is a residential condominium …"] is the record of the POI "Trump World Tower".
2.1.2 Text Data
The text data provides rich contextual knowledge of the city space from different perspectives (e.g., the spatial context) [13], and it plays an important role in geospatial understanding. In this work, we collect two types of text corpus.
Review Data. The review of urban places provides commercial information that citizens use to make business decisions [29], playing a critical role in urban knowledge distillation. We collect review data from Google Map  [^6] https://www.google.com/maps. Specifically, we first manually split the city into multiple rectangular regions, then we utilize the Google Map API to query the places contained within each region and their reviews. Each review record contains a place name, a coordinate geometry value, a rating, and a text review. For example, ["Lifestyles Academy Inc", coordinate (-87.87 41.65), 4.9, "Very nice organization and …"] is the review record of place "Lifestyles Academy Inc".
Web Page Data. The web page data works as the general text corpus for the city, and it contains rich geoscience knowledge that has been utilized in recent urban entity and relation extraction studies [13]. We collect web page data from the Google search engine. Specifically, we first input the name of the crawled AOI, Road, and POI record into Google. Then we concatenate the textual sentences of the top 10 retrieved web pages. Each web page record contains a long urban text description.
2.2 Data Preprocessing
Before constructing the UrbanKGC dataset, we first preprocess the raw datasets. We filter out AOIs, roads, POIs, reviews, and web pages whose crawled textual descriptions are null value, too short (e.g., less than ten word description) or meaningless (e.g., just repeating the POI name). In addition, we remove irrelevant information from the text description, such as non-English characters, non-ASCII gibberish, website addresses, and so on. More details can be found in Appendix A.
3 Preliminary
@@
3.1 Task Definition and Problem Formulation
Before diving into the technical details, we first introduce the definition of UrbanKG:
Definition 1
@@ 3.1 Task Definition and Problem Formulation
The UrbanKG encodes diverse urban semantic knowledge by connecting urban entities into a multi-relational graph. This work aims to construct an UrbanKG from collected unstructured text data. We decompose the UrbanKG construction (UrbanKGC) process into two sequential knowledge graph construction tasks, namely relational triple extraction [10] and knowledge graph completion [11]. We first provide the basic definition for these two subtasks, and then introduce the problem formulation of this work.
3.1.1 Task Definition